OpenStudy (anonymous):

I NEED HELP NOW PLEASE! DESPERATE! the equation of a comic section is 2y^2 - x - 8y + 5 = 0. What is the equation in graphing form.

OpenStudy (anonymous):

if u give me 2 seconds i can get the answer

OpenStudy (anonymous):

Okay I have 15 min left

OpenStudy (anonymous):

\[2y^2-8y=x-5\] \[2\left( y^2-4y+4-4 \right)=x-5\] \[2\left( y-2 \right)^2-8=x-5\] \[2\left( y-2 \right)^2-8=x-5\] \[\left( y-2 \right)^2=\frac{ 1 }{ 2 }\left( x+3 \right)\] it is a parabola

OpenStudy (anonymous):

y-2=0,y=2 x+3=0,x=-3 so vertex is (-3,2)

OpenStudy (anonymous):

The answer choices are A: x - 1 = 2(y - 4)^2 B: x + 3 = 2(y - 2)^2 C: x^2/2 - (y + 5)^2/4 = 1 D: x^2/2 - (y + 5)^2/4= 1

OpenStudy (anonymous):

@barbdewysveck1

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

delete the last line you get the answer.

OpenStudy (anonymous):

\[2\left( y-2 \right)^2=x+3\]

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

correct.